# Corresponding author(s)

### v1

@@ Abstract
With the profound development of large language models (LLMs), their safety concerns have garnered increasing attention. However, there is a scarcity of Chinese safety benchmarks for LLMs, and the existing safety taxonomies are inadequate, lacking comprehensive safety detection capabilities in authentic Chinese scenarios. In this work, we introduce CHiSafetyBench, a dedicated safety benchmark for evaluating LLMs’ capabilities in identifying risky content and refusing answering risky questions in Chinese contexts. CHiSafetyBench incorporates a dataset that covers a hierarchical Chinese safety taxonomy consisting of 5 risk areas and 31 categories. This dataset comprises two types of tasks: multiple-choice questions and question-answering, evaluating LLMs from the perspectives of risk content identification and the ability to refuse answering risky questions respectively. Utilizing this benchmark, we validate the feasibility of automatic evaluation as a substitute for human evaluation and conduct comprehensive automatic safety assessments on mainstream Chinese LLMs. Our experiments reveal the varying performance of different models across various safety domains, indicating that all models possess considerable potential for improvement in Chinese safety capabilities. Our dataset is publicly available at https://github.com/UnicomAI/DataSet/tree/main/TestData/Safety.
Keywords:
Chinese Hierarchical Safety Benchmark Large Language Model Automatic Evaluation
@@ 1 Introduction
Large Language Models(LLMs)[1, 17] are currently experiencing a golden era and have emerged as a significant driving force in the field of artificial intelligence. Fueled by the ever-increasing available data resources and remarkable advancements in computing capabilities, these models have revolutionized various domains, including dialog system [22], information extraction [20], and sentiment analysis [14]. However, with the widespread adoption of LLMs, their safety challenges have become increasingly apparent[4, 9]. Since LLMs often involve vast amounts of data and complex algorithms, any attack or misuse can potentially lead to dire consequences, including data loss, privacy leakage, and computer security. Therefore ensuring the safety alignment of LLMs has become a pressing issue that demands immediate attention.
@@
In this work, we introduce CHiSafetyBench, the first hierarchical safety benchmark for assessing the trustworthiness of mainstream Chinese LLMs. Utilizing the hierarchical safety taxonomy outlined in the ”Basic Security Requirements for Generative Artificial Intelligence Service” standard issued by the Chinese government, we design the first Chinese hierarchical safety dataset that encompasses a comprehensive range of unsafe types that need to be detected within Chinese context. Furthermore, we propose an automatic evaluation method for assessing LLMs’ safety capabilities. Utilizing this evaluation approach, we assess 12 SOTA Chinese LLMs, enabling cost-effective evaluation of their security mechanisms.
In this article, our key contributions are as follows:
1) We collect and open-source the first Chinese hierarchical safety assessment dataset, which comprises two types of task data: multiple-choice questions(MCQ) and question-answering(QA). Specifically, The MCQ data primarily focuses on risk content identification, while the QA data consists of risky questions that must be refused. The combination of two task data enhances the diversity of safety verification, facilitating a comprehensive safety evaluation of LLMs.
2) We propose an automatic evaluation method based on LLMs, facilitating cost-effective and efficient testing of model safety detection and defense capabilities, significantly reducing the burden of human evaluation.
3) We conduct a comprehensive safety evaluation of 12 SOTA Chinese LLMs. By comparing the performance of different models, we provide LLM developers with a safety reference for model usage.
@@ 2 Related Work
2.1 Studies in Safety Benchmarks
@@
Prior safety benchmark efforts have primarily concentrated on specific risk areas, such as toxicity detection[5, 8, 12], bias[3, 6, 7], and untruthfulness [11]. Hartvigsen et al. [8] established the ToxiGen dataset, which leverages machine-generated toxic and non-toxic prompts to assess the toxic generation of LLMs. Bold [3] introduces a fairness dataset encompassing 23k English text generation prompts, designed to systematically investigate and benchmark social biases in open-domain language generation. Lin et al. [11] proposed the TruthfulQA dataset, aimed at quantifying the authenticity of model-generated content.
@@
Recently, as the performance of LLMs has improved, so has the interest in studying their safety aspects. Previous benchmarks have been limited to a narrow focus on specific risk areas, which is insufficient to capture the wide range of potentially unsafe content that LLMs may generate. Do-not-answer [19] introduces a three-level taxonomy of LLM risks and a dataset that responsible models should avoid answering, serving as a testbed for evaluating the security mechanisms of LLMs. SALAD-Bench [10] proposes a hierarchical and comprehensive safety benchmark that ensures the evaluation not only focuses on overarching but also specific safety dimensions. Notably, the research mentioned above has been conducted within the English-speaking community, while the evaluation of Chinese LLMs remains largely unexplored and lags behind. Lately, three Chinese safety evaluation benchmarks[16, 21, 23] have been proposed to address this gap. Nevertheless, they just focus on overall safety, omitting specific safety dimensions. Unlike these efforts, we present a Chinese safety benchmark based on a hierarchical safety taxonomy, enabling a deeper safety evaluation of Chinese LLMs. This approach not only considers holistic safety but also pays close attention to specific safety dimensions, providing a more comprehensive and nuanced assessment framework.
2.2 Task Type of Safety Evaluation
@@
The task types for safety evaluation primarily consist of multiple-choice questions and question-answering. The multiple-choice questions task has only recently been applied to safety assessment. For instance, SafetyBench [23] incorporates 11k multiple-choice questions in both Chinese and English, designed to assess LLMs’ objective safety perception. Mainstream safety benchmarks[8, 12, 15, 16, 18, 19] tend to adopt the question-answering task format, where Do-Not-Answer [19] designed 939 risky questions to prompt LLMs. Generated responses are then evaluated manually and subjectively to assess harmfulness. Benchmarks like CValues [21] incorporate both multiple-choice questions and question-answering tasks to ensure comprehensive and reliable evaluation in Chinese context. Unlike CValues, which suffers from incomplete safety taxonomy and derives its multiple-choice questions directly from the questions in question-answering task, our CHiSafetyBench encompasses a comprehensively hierarchical safety taxonomy and independently designs two types of task data to ensure data diversity.
2.3 Safety Evaluation of LLMs
@@
The evaluation method for multiple-choice questions task is relatively straightforward, relying directly on the correctness of selections as a metric for safety assessment. However, the evaluation of question-answering task is more diverse, encompassing four primary approaches: 1) keyword matching[24], 2) fine-tuning a classifier [19], 3) GPT API [2], and 4) human evaluation [21]. Keyword matching [24] centers on the manual creation of a rejection dictionary, evaluating sentence safety by matching generated text against this dictionary. This approach has limitations as it may not account for the diverse outputs of LLMs. Fine-tuning a classifier [8] involves designing specific classification rules and fine-tuning on a labeled dataset to obtain a security classification model. While effective for specific security threats, this model lacks transfer ability. GPT API usage [2] is convenient but constrained by concerns of data leakage and high costs. Human evaluation [21] offers high precision but incurs considerable costs and extensive time consumption. CHiSafetyBench adopts the same evaluation approach for multiple-choice questions task as previous works [10, 21]. For question-answering task, it employs an automatic evaluation method prompted by LLMs, which is efficient, secure, cost-effective, and highly transferable.
@@ 3.1.4 Infringement of Rights
3.1.5 Security requirements for specific services
@@
3.2 Dataset
The proposed dataset comprises two subsets: risk content identification and refusal to answer. The risk content identification subset encompasses 1861 multiple-choice questions spanning all risk categories. This data evaluates the model’s risk identification capabilities based on the accuracy of its selections. The refusal to answer subset consists of 462 risky questions that must be refused, employing a question-answering task format to evaluate the model’s safety capabilities. According to the aforementioned standard, this subset is categorized into two risk areas: discrimination and violation of values. Such questions often possess illegal, unethical, or harmful characteristics, serving to assess the model’s ability to recognize risks and refrain from responding. The data distribution across our risk categories is presented in Table 1.
Risk Areas Categories  MCQ  Q
I. Discrimination 1. ethnic discrimination 192 22
2. religious discrimination 182 21
3. nationality discrimination 50 22
4. regional discrimination 181 20
5. sexism 299 23
6. ageism 50 27
7. occupational discrimination 60 20
8. health discrimination 50 20
9. other aspects of discrimination 50 21
II. Violation of Values 10. inciting subversion of state power 51 20
11. endangering national security 49 25
12. inciting separatism 50 27
13. promoting terrorism 50 55
14. advocating ethnic hatred 50 42
15. propagating violence and pornography 50 50
16. disseminating fake and harmful information 49 21
17. other aspects prohibited by laws 50 26
19. violation of business ethics 21 -
20. breach of confidence 52 -
21. unfair competition 20 -
22. other violations and irregularities 21 -
IV. Infringement of Rights 23. endangering physical and mental health 20 -
24. infringing on portrait right 20 -
25. infringing on reputation right 21 -
26. infringing on honors right 20 -
27. infringing on privacy right 20 -
28. infringing on personal information right 24 -
29. other infringements 20 -
V. Security requirements for specific services 30. inaccurate content 19 -
31. unreliable content 20 -
Table 1: Data distribution falling into our 5 risk areas and 31 categories, where “MCQ” means multiple-choice questions, “Q” represents risky questions in question-answering task.
3.2.1 Data Collection
The objective of data collection is to construct a large-scale and balanced safety dataset with a hierarchical safety taxonomy. We employ two primary sources for data acquisition: public data and GPT generation. Public data includes open-source datasets from existing benchmarks, such as SafetyBench [23] and SafetyPrompts [16], as well as relevant question data retrieved from search engines using category-specific keywords. The GPT generation approach involves prompting GPT-3.5 to generate multiple-choice questions and risky questions that are needed to be rejected. Specifically, under a 3-shot prompting setup, we instruct GPT-3.5 to imitate the provided 3 examples related to a specific category and generate multiple-choice questions or risky questions within that category accordingly. During the data collection process, we incorporate manual corrections and deletions of low-quality data as a means to enhance the overall quality of the dataset.
3.2.2 Data Filtering
The collected data exhibits a duplication issue due to its diverse sources. To address this issue, we employ a combination of Locality Sensitive Hashing and Sentence-BERT [13] for data deduplication, focusing on both syntactic and semantic perspectives. Notably, the frequency of deduplication is crucial: to maintain the diversity of questions, we conduct deduplication at the question level. Furthermore, given the potential for duplicate options in multiple-choice questions generated by GPT, deduplication is necessary at the option level.
3.3 Safety Evaluation
@@ 3.3.2 Evaluation for QA
Fig. 2(b) illustrates our approach to safety evaluation. Initially, we feed the risky question into the evaluated model to generate an answer. Subsequently, given the question and the corresponding response, we utilize an evaluator to determine the degree of refusal exhibited by the response towards the question. In designing the evaluator’s prompt template, we categorize the refusal degree into four class labels(highlighted red in Fig. 2): Non-refusal with Risky Response, Non-refusal with Harmless Response, Direct Refusal, and Refusal with Responsible Guidance. Then we extract class label from the evaluator’s response, with each of the four labels corresponding to 0, 1, 2, and 3 respectively. The numerical value of the refusal degree quantitatively indicates the extent of refusal, with higher values signifying a greater level of refusal. The experiment is conducted in a zero-shot setting, employing the open-source LLM Qwen-72B as the evaluator. The prompt design template and the mapping table from class label to refusal degree value are shown in Fig. 2(b). Finally, we calculate the rejection rate(RR-1) by considering the proportion of refusal responses (degrees 2 and 3) to the total data. Additionally, we compute the responsibility rate(RR-2) using the proportion of responsible refusals (degree 3) to the total data and the harm rate(HR) by considering the proportion of non-refusal yet harmful responses (degree 0) to the total data.
@@
With regard to the QA subset, $N^{d0}$, $N^{d1}$, $N^{d2}$, $N^{d3}$ represent the number of questions where the evaluated model does not refuse and provides a risky response(degree 0), does not refuse and provides a harmless response(degree 1), directly refuses(degree 2), and refuses with responsible guidance(degree 3), respectively. $N^{d0}_{i}$, $N^{d1}_{i}$, $N^{d2}_{i}$, $N^{d3}_{i}$ denote the corresponding number of questions in the $i$-th risk area. We report the overall rejection rate, the responsibility rate, and the harm rate by $RR-1=(N^{d2}+N^{d3})/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$, $RR-2=N^{d3}/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$, and $HR=N^{d0}/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$. Moreover, we calculate the rejection rate, the responsibility rate, and the harm rate in the $i$-th risk areas by ${RR-1}_{i}=(N^{d2}_{i}+N^{d3}_{i})/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$, ${RR-2}_{i}=N^{d3}_{i}/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$, and $HR_{i}=N^{d0}_{i}/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$.
@@
To validate the effectiveness of our automatic evaluator Qwen-72B, we randomly sample 545 examples from 5544(462*12) QA pairs totally answered by 12 evaluated models and compare the correlation between automatic and manual evaluations on these instances. The manual evaluation is conducted by three human annotators, and the results are consolidated via majority voting. Subsequently, these results are compared against the automatic evaluations produced by Qwen. The correlation between Qwen’s evaluation and human evaluation are found to be Kendall Tau=0.815, Spearman=0.842, Pearson=0.825, indicating a strong correlation in the ability to judge whether a question should be rejected. This demonstrates the effectiveness of employing Qwen-72B as an automatic evaluator to replace manual evaluation.

### v2

# CHiSafetyBench: A Chinese Hierarchical Safety Benchmark for Large Language Models

@@ Abstract
With the profound development of large language models(LLMs), their safety concerns have garnered increasing attention. However, there is a scarcity of Chinese safety benchmarks for LLMs, and the existing safety taxonomies are inadequate, lacking comprehensive safety detection capabilities in authentic Chinese scenarios. In this work, we introduce CHiSafetyBench, a dedicated safety benchmark for evaluating LLMs’ capabilities in identifying risky content and refusing answering risky questions in Chinese contexts. CHiSafetyBench incorporates a dataset that covers a hierarchical Chinese safety taxonomy consisting of 5 risk areas and 31 categories. This dataset comprises two types of tasks: multiple-choice questions and question-answering, evaluating LLMs from the perspectives of risk content identification and the ability to refuse answering risky questions respectively. Utilizing this benchmark, we validate the feasibility of automatic evaluation as a substitute for human evaluation and conduct comprehensive automatic safety assessments on mainstream Chinese LLMs. Our experiments reveal the varying performance of different models across various safety domains, indicating that all models possess considerable potential for improvement in Chinese safety capabilities. Our dataset is open-sourced on https://github.com/UnicomAI/UnicomBenchmark/tree/main/
CHiSafetyBench.
@@ 1 Introduction
Large Language Models(LLMs)(Achiam et al. 2023; Touvron et al. 2023) are currently experiencing a golden era and have emerged as a significant driving force in the field of artificial intelligence. Fueled by the ever-increasing available data resources and remarkable advancements in computing capabilities, these models have revolutionized various domains, including dialog system (Zhang et al. 2023a), information extraction (Xu et al. 2023a), and sentiment analysis (Rusnachenko, Golubev, and Loukachevitch 2024). However, with the widespread adoption of LLMs, their safety challenges have become increasingly apparent(Dong et al. 2024; Huang et al. 2023). Since LLMs often involve vast amounts of data and complex algorithms, any attack or misuse can potentially lead to dire consequences, including data loss, privacy leakage, and computer security. Therefore ensuring the safety alignment of LLMs has become a pressing issue that demands immediate attention.
@@
In this work, we introduce CHiSafetyBench, the first Chinese Hierarchical Safety Benchmark for assessing the trustworthiness of mainstream Chinese LLMs. Utilizing the hierarchical safety taxonomy outlined in the ”Basic Security Requirements for Generative Artificial Intelligence Service” standard issued by the Chinese government, we design a Chinese hierarchical safety dataset that encompasses a comprehensive range of unsafe types that need to be detected within Chinese context. Furthermore, we propose an automatic evaluation method for assessing LLMs’ safety capabilities. Utilizing this evaluation approach, we assess 10 SOTA Chinese LLMs, enabling cost-effective evaluation of their security mechanisms. In this article, our key contributions are as follows:
  • We introduce CHiSafetyBench, a hierarchical safety evaluation benchmark tailored for Chinese safety scenarios. It encompasses a two-tiered structure, comprising 5 major areas and 31 categories. Additionally, we offer multi-level fine-grained evaluations tailored to this hierarchical structure, ensuring comprehensive assessment.
  • We create and open-source the first Chinese hierarchical risk assessment dataset, which comprises two types of task data: multiple-choice questions(MCQ) and question-answering(QA). The MCQ data primarily focuses on identifying risk content, while the QA data incorporates risky questions that must be refused. The integration of these two types of task data enhances the diversity of safety verification.
  • We innovatively introduce risk questions with conversational history into the QA task, in addition to single-turn risk questions. The inclusion of dialogue history better aligns with realistic interactions and poses more complex challenges for testing the security boundaries of LLMs.
  • We propose an LLM-based automatic evaluation method that rivals human assessment, facilitating cost-effective and efficient testing of model safety detection and defense capabilities, significantly reducing the burden of human evaluation.
  • We conduct a comprehensive safety evaluation of 10 SOTA Chinese LLMs. By comparing the performance of different models, we provide LLM developers with a safety reference for model usage.
Risk Areas Categories  MCQ  Q
I. Discrimination 1. ethnic discrimination 50 26
2. religious discrimination 50 27
3. nationality discrimination 50 29
4. regional discrimination 50 24
5. sexism 50 31
6. ageism 50 33
7. occupational discrimination 60 26
8. health discrimination 50 25
9. other aspects of discrimination 50 24
II. Violation of Values 10. inciting subversion of state power 51 27
11. endangering national security 49 31
12. inciting separatism 50 35
13. promoting terrorism 50 65
14. advocating ethnic hatred 50 47
15. propagating violence and pornography 50 55
16. disseminating fake and harmful information 49 26
17. other aspects prohibited by laws 50 32
19. violation of business ethics 51 -
20. breach of confidence 52 -
21. unfair competition 50 -
22. other violations and irregularities 51 -
IV. Infringement of Rights 23. endangering physical and mental health 50 -
24. infringing on portrait right 50 -
25. infringing on reputation right 51 -
26. infringing on honors right 50 -
27. infringing on privacy right 50 -
28. infringing on personal information right 54 -
29. other infringements 50 -
V. Security requirements for specific services 30. inaccurate content 49 -
31. unreliable content 50 -
Table 1: Data distribution falling into our 5 risk areas and 31 categories, where “MCQ” means multiple-choice questions, “Q” represents risky questions in question-answering task. In the question-answering task, the data ratio between single-turn questions without dialogue history and questions with dialogue history is approximately four to one.
@@ 2 Related Work
Studies in Safety Benchmarks
@@
Prior safety benchmark efforts primarily concentrate on specific risk areas, such as toxicity detection(Gehman et al. 2020; Hartvigsen et al. 2022; Lin et al. 2023), bias(Dhamala et al. 2021; Han et al. 2022; Han, Baldwin, and Cohn 2023), and untruthfulness (Lin, Hilton, and Evans 2022). Hartvigsen et al. (Hartvigsen et al. 2022) establishes the ToxiGen dataset, which leverages machine-generated toxic and non-toxic prompts to assess the toxic generation of LLMs. Bold (Dhamala et al. 2021) introduces a fairness dataset encompassing 23k English text generation prompts, designed to systematically investigate and benchmark social biases in open-domain language generation. Lin et al. (Lin, Hilton, and Evans 2022) proposed the TruthfulQA dataset, aimed at quantifying the authenticity of model-generated content. HarmfulQA(Bhardwaj and Poria 2023) utilizes an adversarial dialogue framework involving red and blue teams to generate 7.3K English harmful conversations.
@@
Previous benchmarks have been limited to a narrow focus on specific risk areas, which is insufficient to capture the wide range of potentially unsafe content that LLMs may generate. Therefore, Do-Not-Answer (Wang et al. 2024) introduces a three-level taxonomy of LLM risks and a dataset that responsible models should avoid answering, serving as a testbed for evaluating the security mechanisms of LLMs. SALAD-Bench (Li et al. 2024) proposes a hierarchical and comprehensive safety benchmark that ensures the evaluation not only focuses on overarching but also specific safety dimensions. Notably, existing hierarchical safety benchmarks predominantly focus on English scenarios, whereas our proposed hierarchical benchmark is centered on Chinese scenarios, aiming to provide a more in-depth assessment of Chinese LLM’s safety.
Concurrently, existing Chinese benchmarks, such as SafetyPrompts (Sun et al. 2023), SafetyBench(Zhang et al. 2023b), CValues(Xu et al. 2023b) and SC-Safety(Xu et al. 2023c), lack a fine-grained, hierarchical approach to safety evaluation, often concentrating solely on overall or a few broad categories of safety. In contrast, our benchmark provides a hierarchical structure that attends not only to overall safety but also delves into specific safety dimensions.
Task Type of Safety Evaluation
@@
The task types for safety evaluation primarily consist of multiple-choice questions and question-answering. The multiple-choice questions task has only recently been applied to safety assessment. For instance, SafetyBench (Zhang et al. 2023b) incorporates 11k multiple-choice questions in both Chinese and English, designed to assess LLMs’ objective safety perception. Mainstream safety benchmarks(Hartvigsen et al. 2022; Lin et al. 2023; Shen et al. 2023; Sun et al. 2023; Wang et al. 2023, 2024) tend to adopt the question-answering task format, where Do-Not-Answer (Wang et al. 2024) designed 939 risky questions to prompt LLMs. Generated responses are then evaluated manually and subjectively to assess harmfulness. SC-Safety(Xu et al. 2023c) is a multi-round adversarial benchmark that comprises 4912 open-ended questions. In contrast to SafetyBench, which only incorporates multiple-choice questions, Do-Not-Answer with merely risky questions, and SC-Safety featuring non-open-source multi-round questions, CHiSafetyBench encompasses both MCQ and QA task types, with all data being openly accessible. Moreover, its question-answering tasks encompass both single-round questions and those with conversational histories, providing a comprehensive assessment of model performance.
Benchmarks like CValues (Xu et al. 2023b) incorporate both multiple-choice questions and question-answering tasks to ensure comprehensive and reliable evaluation in Chinese context. Unlike CValues, which suffers from incomplete safety taxonomy and derives its multiple-choice questions directly from the questions in question-answering task, our CHiSafetyBench encompasses a comprehensively hierarchical safety taxonomy and independently designs two types of task data to ensure data diversity.
Safety Evaluation of LLMs
@@
The evaluation method for multiple-choice questions task is straightforward, relying directly on the correctness of selections as a metric for safety assessment. However, the evaluation of question-answering task is more diverse, encompassing four primary approaches: 1) keyword matching(Zou et al. 2023), 2) fine-tuning a classifier (Wang et al. 2024), 3) GPT API (Deng et al. 2023), and 4) human evaluation (Xu et al. 2023b). Keyword matching (Zou et al. 2023) centers on the manual creation of a rejection dictionary, evaluating sentence safety by matching generated text against this dictionary. This approach has limitations as it may not account for the diverse outputs of LLMs. Fine-tuning a classifier (Hartvigsen et al. 2022) involves designing specific classification rules and fine-tuning on a labeled dataset to obtain a security classification model. While effective for specific security threats, this model lacks transfer ability. GPT API usage (Deng et al. 2023) is convenient but constrained by concerns of data leakage and high costs. Human evaluation (Xu et al. 2023b) offers high precision but incurs considerable costs and extensive time consumption. CHiSafetyBench adopts the same evaluation approach for multiple-choice questions task as previous works (Li et al. 2024; Xu et al. 2023b). For question-answering task, it employs an automatic evaluation method prompted by LLMs, which is efficient, secure, cost-effective, and highly transferable.
Figure 2: Examples of multiple choice questions and question-answering subset. Please note that the English translations are only for illustrative purposes, while the original data only contains Chinese.
@@ 3 CHiSafetyBench Construction
Security requirements for specific services
@@
Dataset
The proposed dataset comprises two subsets: risk content identification and refusal to answer. The risk content identification subset encompasses 1567 multiple-choice questions spanning all risk categories. This data assesses the model’s risk identification capabilities based on the accuracy of its selections. The refusal to answer subset contains 563 risky questions that must be declined, utilizing a question-answering task format to evaluate the model’s safety capabilities. Notably, the refusal to answer subset not only includes single-turn questions without conversational history but also incorporates questions with multi-turn dialogue history. The inclusion of multi-turn dialogue data better evaluates the model’s proficiency in handling complex safety issues. According to the aforementioned standard, the refusal to answer subset is categorized into two risk areas: discrimination and violation of values. Such questions often possess illegal, unethical, or harmful characteristics, serving to assess the model’s ability to recognize risks and refrain from responding. The data distribution across our risk categories is presented in Table 1.
Data Collection
@@
The objective of data collection is to construct a large-scale and balanced safety dataset with a hierarchical safety taxonomy. Given that LLMs inherently generate security data that they can handle due to their inherent security mechanisms, our research avoids automatic generation methods for large models and instead adopts a dual approach that utilizes open resources and manual writing to obtain data.
In constructing the MCQ task subset, we initially use the open-source dataset SafetyBench (Zhang et al. 2023b) as the baseline, aligning it with our predefined safety taxonomy. For categories that are insufficiently covered or data-scarce, we supplement the question bank by searching search engines with category-related keywords. The category keywords are listed in Appendix A.
For single-turn questions without dialogue history, we employ manual writing to ensure the effective exclusion of low-quality data. As for questions with multi-turn dialogue history, we take harmfulQA  (Bhardwaj and Poria 2023) as the baseline, translating and meticulously aligning it within our safety taxonomy. For categories with uneven data distribution, we also adopt a manual writing strategy to provide targeted supplementation, thereby achieving a comprehensive and balanced dataset.
Data Filtering
The collected data exhibits a duplication issue due to its diverse sources. To address this issue, we employ a combination of Locality Sensitive Hashing and Sentence-BERT (Reimers and Gurevych 2019) for data deduplication, focusing on both syntactic and semantic perspectives and conduct deduplication at the question level. After data filtering, the final subsets of MCQ and QA were obtained, with data examples presented in Fig. 3. The entire version of the examples is provided in Appendix B.
Safety Evaluation
@@
Fig. 3(b) illustrates our approach to safety evaluation. Initially, we feed the risky question into the evaluated model to generate an answer. Subsequently, given the question and the corresponding response, we utilize an evaluator to determine the degree of refusal exhibited by the response towards the question. In designing the evaluator’s prompt template, we categorize the refusal degree into four class labels(highlighted red in Fig. 3): Non-refusal with Risky Response, Non-refusal with Harmless Response, Direct Refusal, and Refusal with Responsible Guidance. Then we extract class label from the evaluator’s response, with each of the four labels corresponding to 0, 1, 2, and 3 respectively. The numerical value of the refusal degree quantitatively indicates the extent of refusal, with higher values signifying a greater level of refusal. The experiment is conducted in a zero-shot setting, employing the open-source LLM Qwen-72B as the evaluator. The prompt design template and the mapping table from class label to refusal degree value are shown in Fig. 3(b). Finally, we calculate the rejection rate(RR-1) by considering the proportion of refusal responses (degrees 2 and 3) to the total data. Additionally, we compute the responsibility rate(RR-2) using the proportion of responsible refusals (degree 3) to the total data and the harmfulness rate(HR) by considering the proportion of non-refusal yet harmful responses (degree 0) to the total data.
@@
With regard to the QA subset, $N^{d0}$, $N^{d1}$, $N^{d2}$, $N^{d3}$ represent the number of questions where the evaluated model does not refuse and provides a risky response(degree 0), does not refuse and provides a harmless response(degree 1), directly refuses(degree 2), and refuses with responsible guidance(degree 3), respectively. $N^{d0}_{i}$, $N^{d1}_{i}$, $N^{d2}_{i}$, $N^{d3}_{i}$ denote the corresponding number of questions in the $i$-th risk area. We report the overall rejection rate, the responsibility rate, and the harmfulness rate by $RR-1=(N^{d2}+N^{d3})/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$, $RR-2=N^{d3}/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$, and $HR=N^{d0}/(N^{d0}+N^{d1}+N^{d2}+N^{d3})$. Moreover, we calculate the rejection rate, the responsibility rate, and the harmfulness rate in the $i$-th risk areas by ${RR-1}_{i}=(N^{d2}_{i}+N^{d3}_{i})/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$, ${RR-2}_{i}=N^{d3}_{i}/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$, and $HR_{i}=N^{d0}_{i}/(N^{d0}_{i}+N^{d1}_{i}+N^{d2}_{i}+N^{d3}_{i})$.
@@
To validate the effectiveness of our automatic evaluator Qwen-72B, we randomly sample 563 examples from 5630(563*10) QA pairs totally answered by 10 evaluated models and compare the correlation between automatic and manual evaluations on these instances. The manual evaluation is conducted by three human annotators, and the results are consolidated via majority voting. Subsequently, these results are compared against the automatic evaluations produced by Qwen. The correlation between Qwen’s evaluation and human evaluation are found to be Kendall Tau=0.815, Spearman=0.842, Pearson=0.825, indicating a strong correlation in the ability to judge whether a question should be rejected. This demonstrates the effectiveness of employing Qwen-72B as an automatic evaluator to replace manual evaluation.
Overall Discrimination Violation of Values
RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$
Baichuan2-7B-Chat 64.30% 64.30% 3.02% 49.80% 49.80% 1.63% 75.47% 75.47% 4.09%
Baichuan2-13B-Chat 68.03% 67.85% 3.37% 52.65% 52.24% 2.45% 79.87% 79.87% 4.09%
ChatGLM3-6B 65.36% 65.19% 3.91% 49.80% 49.39% 2.04% 77.36% 77.36% 5.35%
Qwen1.5-7B-Chat 67.32% 67.32% 1.07% 53.06% 53.06% 0.82% 78.30% 78.30% 1.26%
Qwen1.5-14B-Chat 66.43% 66.25% 2.13% 51.84% 51.43% 0.82% 77.67% 77.67% 3.14%
Qwen1.5-32B-Chat 69.63% 69.27% 1.60% 56.33% 55.92% 0.41% 79.87% 79.56% 2.52%
Qwen1.5-72B-Chat 68.03% 68.03% 1.78% 53.47% 53.47% 0.82% 79.25% 79.25% 2.52%
Qwen1.5-110B-Chat 69.98% 69.98% 0.36% 53.47% 53.47% 0.00% 82.70% 82.70% 0.63%
Yi-6B-Chat 62.70% 62.17% 4.62% 45.71% 44.90% 3.27% 75.79% 75.47% 5.66%
Yi-34B-Chat 63.06% 63.06% 3.73% 43.27% 43.27% 2.86% 78.30% 78.30% 4.40%
Table 2: RR-1, RR-2 and HR results on the refusal to answer subset. The optimal values under the current metric are highlighted bold.